Alto Company issued 1000 shares of 7% preferred stock with a $100 par value. This means that:

Multiple Choice

  • The maximum total dividends preferred stockholders have a right to for any given year is $7

  • The maximum total dividends preferred stockholders have a right to for any given year is $70

  • Preferred shareholders have a guaranteed dividend.

  • The maximum total dividends preferred stockholders have a right to for any given year is $7000
Answer #1

Dividend = 100 * 7 % = $ 7 per preferred stock

This dividend is needed to be paid only if there is income available.

maximum total dividends preferred stockholders =  1000 * 100 * 7 % = $ 7000

OPTION D IS CORRECT.